Calculate the magnitude of the gravitational force of attraction between two spherical of lead with a mass of 10 kg and 6 kg respectively if are placed 50 mm apart
1
Expert's answer
2021-08-17T16:13:15-0400
"F=G\frac{mM}{r^2}\\F=6.67\cdot10^{-11}\frac{10\cdot6}{0.05^2}=1.6\cdot10^{-6}N"